# Context-Aware Code Generation Framework for Code Repositories: Local, Global, and Third-Party Library Awareness

### 18 Dec 2023 v2

@@ 3.1.2. Third-Party Library Extraction
As third-party libraries commonly appear in code files, it is often the case that when we provide specific third-party libraries to a model, the model is likely to have already learned how to use those libraries during the training process [45, 46]. Therefore, we only need to provide the name of the library and the model will be able to automatically understand how to use that library to perform a specific task. For example, we can ask the model to use the NumPy library to compute the cosine similarity, given the input vector_a and vector_b. In this case, the model can directly call functions in the NumPy library, such as numpy.dot and numpy.linalg.norm, to perform the required computations. The generated code example is as follows: $cosine\_similarity=numpy.dot(vector\_a,vector\_b)/(np.linalg.norm(vector\_a)*numpy.linalg.norm(vector\_b)).$ By simply providing the name of the library and the task requirements, the model can generate the appropriate code, helping the developer accomplish programming tasks quickly. The pre-trained model’s prior knowledge and extensive large-scale training enable them to effectively utilize third-party libraries, improving the efficiency and accuracy of code generation. Therefore, we extracted the import statements of third-party libraries directly from all code files via AST and saved their library names to construct the third-party library base.
3.2. Three Types of Repo-Aware Knowledge Retrieval
@@
As we discussed in Section 2.1, although the maximum input length grows greatly, it is still impossible to feed the entire code repo and third-party library information to LLMs. Furthermore, irrelevant context in prompts not only affects model efficiency and increases expenses, but also potentially compromises the model performance [33].
@@
To enable the model to fully utilize the three types of information extracted from the code repository during code generation, we conduct comprehensive knowledge mining: extracting four types of local information (Section 3.2.1), retrieving global functions from the function base using developer requirements to obtain global-aware information (Section 3.2.2), and structuring third-party library information from the third-party library base (Section 3.2.3).
Figure 4. Local-Module-Aware, Global-Module-Aware, Third-Party-Library-Aware Knowledge
3.2.1. Local-Module Knowledge Mining
@@
  • Local functions. Local functions encompass all functions that are defined and implemented within the local module. We extract these local functions and represent them in the form of ¡Fully Qualified Name (FQN), Function Summary, Function Signature¿. The FQN provides details about the function’s location within the module, the Function Summary reflects the natural language description of the function’s purpose, and the Function Signature reveals its input and output specifications. Fig. 4 (a) illustrates local functions, displaying the FQN, summary, and signature for each function, such as function k. We use indentation to differentiate whether a function belongs to a class. For example, function 1 belongs to the HTMLDocument class, indicated by an indented tab, while function k is a module function not associated with any class. Additionally, each function’s information is enclosed within {} braces. This structuring serves the purpose of organizing the information for all local functions, aiding the model in more effectively recognizing local function knowledge.
  • Class Instance Attributes. Class instance attributes, also known as instance variables, are variables defined within a class’s constructor function (__init__ in Python). These variables are specific to instances of the class and store instance-specific attributes or state information. Some local functions may be defined within a class, and they might involve interactions with class instance attributes, accessing or modifying them to achieve specific functionality. For instance, in Fig. 4 (a), the source code within function1 explains how class instance attributes are defined. Therefore, to represent class attributes, we do not extract them separately but include the source code of the constructor function (__init__) that defines these class attributes directly.
  • Local Module FQN. Local Module FQN partially reflects the structural information of the working context (local module) within the entire code repository. It helps the model understand the context of how local functions, class instance attributes, and variables within this module are used. We derive the Local Module FQN from the current file path, providing a contextual reference for the code within this module. For example, the local module FQN of the file in Fig. 4 (b) is unstructured.documents.html.
@@
After obtaining these four types of local-module knowledge, we organize this knowledge in a structured format. For instace, as illustrated in Fig. 4 Local Module Aware Knowledge.
3.2.2. Global-Module Knowledge Mining
@@
Figure 5. Two Types of Code Generator Units
@@
  • What-If Code Generator Unit : As illustrated in Fig. 5 (a), in the task description, we use a structured prompt that defines roles, task requirements, and output requirements for the LLM. In the prompt, we define two terminologies: developer_requirement, and to_be_generated_code. The developer_requirement contains the function description and function definition. In task description part (highlighted in purple), we set three rules and use commands to guide the model make chain-of-thought. Then, we provide three representative examples for in-context learning. The inputs is the developer_requirement (highlighted in green), and output is the to_be_generated_code (highlighted in red). Notably, the generated code is only used to retrieve the global function, so we do not require its completeness as it is not the final code. This helps the retrieval of relevant reusable functions in the next Global Function Retrieval Unit.
@@
  • Global Function Retrieval Unit: This unit uses function descriptions and the what-if function-code as input to retrieve reusable global functions from the function base. Specifically, we first embed function_description into $vector_{des}$ and compute the cosine similarity between this vector and the summary vectors of all functions in the function base. Thus, we can rank and obtain the top k $RetrievedFun_{des}$. Second, we embed the what-if function-code as a $vector_{f}$ and compute the cosine similarity between this vector and the code vectors of all the functions in the function base, so we can rank and obtain the top k $RetrievedFun_{code}$. Finally, we merge the sets of $RetrievedFun_{des}$ and $RetrievedFun_{d}es$ to obtain the global function that is most relevant to the provided input, which is denoted as $RetrievedFun$
@@
  • Information Unification Unit: Above all, for each $RetrievedFun$, we organize and display the retrieved information as ¡FQN, function signature, summary¿ in a structured format. Fig. 4 (c) presents the global-aware knowledge after unification. Similar to local functions, the information for each function is enclosed within {} braces to assist the model in more efficiently recognizing function knowledge.
3.2.3. Third-Party Library Knowledge Mining
@@
In the process of code generation, to behavior similar to human developers, it’s crucial for the LLM to be aware of the third-party libraries pre-installed in the current code repository environment. Without this awareness, there’s a risk of generating code that relies on third-party libraries unsupported by the current environment, rendering the code non-functional. To ensure that, we create a comprehensive list of all the third-party libraries found in the third-party library base, as extracted in section 3.1.2. It’s important to note that we extract only the names of these libraries and do not include any of their code. Third-party libraries are publicly accessible and widely applied, so according to our empirical observation, we found that the model has already encountered and learned most of the libraries during its training process. Thus, we only extract the names of these libraries without any of their code. By constructing this list of third-party libraries, we empower the LLM to generate code that aligns with the available environment and utilizes the appropriate libraries effectively, reducing the risk of non-executable code.
3.3. $A^{3}$ode Generation
@@
After section 3.2, we obtain structured local-aware knowledge, global-aware knowledge, and third-party library-aware knowledge. To ensure the model’s full awareness of these three types of knowledge, we design an $\mathbf{A^{3}}$ Code Generation Unit that seamlessly integrates user requirements with the three knowledge sources to generate $A^{3}$ functional code.
@@
The $A^{3}$ Code Generation Unit, as depicted in Fig. 5 (b), consists of four key components: task description, examples, input, and output. In the task description, we guide the model through the process of code generation step by step, employing a chain-of-thought approach [47, 48, 49].
@@
  • Additionally, we have included an @command that reiterates the importance of strictly following the steps from 1 to 5.
We establish three rules to govern the code generated by the model, which include calling import FQN, restricting the use of specified third-party libraries, and ensuring that the generated code is not empty (as occasionally, the model may generate abstract functions) We then provide 3 representative examples. Finally, we input our 3 knowledge into Unit to get the $A^{3}$ function code (i.e., _to_be_generated_code in Fig. 5 (b)).
4. Experiment Setting
In this section, we describe our dataset and evaluation metrics used to estimate the performance of our framework. We select GPT- 3.5-Turbo as the default foundation LLM for this study.
We propose the following four Research Questions (RQs):
@@ 4. Experiment Setting
4.1. Dataset Collection
While there exist some benchmark datasets designed for code generation, such as [50, 51, 34, 52, 16], They suffer from the following issues: 1) some of the existing datasets contain code repositories that are already present in the training corpora of popular LLMs, leading to an overestimation of the code repository-aware code generation capabilities. 2) existing benchmarks only evaluate class-level or small unit test cases, which is not suitable for our research goals. Therefore, to address the potential risk of data leakage and accurately reflect the code generation capability, we created a new benchmark dataset, dubbed RepoEval. This decision stems from our intention to verify whether the models truly grasp $A^{3}$ code generation after receiving the information we provide.
4.1.1. Code Repository Collection
According to OpenAI’s official website [^4] https://platform.openai.com/docs/models/gpt-3-5, the training data for GPT-3.5-Turbo is cut off as of September 2021. To ensure that our benchmark dataset contains code repositories that have not been seen during model training, we select projects with the earliest release date after October 1, 2021 on PyPI. We further filter the code repositories by the following criteria: The corresponding GitHub code repositories have more than 100 stars and 100 forks. As the number of stars and branches approximately indicates the quality of the repository. We further selected repositories with more than 10 code files to ensure they have a sufficient number of global modules to help create the benchmark dataset. Ultimately, we collected 29 code repositories that meet these criteria for creating the benchmark.
Then, we constructed 29 function bases and 29 third-party library bases for these 29 code repositories following the method described in Section 3.1. In the function bases extracted for these 29 code repositories, there are a total of 13,784 functions. Among these functions, 10,077 are class member functions, while the remaining 3,707 are standalone functions defined within modules. Since class member functions often involve complex relationships related to inheritance and other classes, it becomes challenging to analyze the code invocation relationships. Therefore, our benchmark dataset in this study focuses solely on standalone functions. From these 3,707 standalone functions, we filter out functions that have comments and are non-empty. In total, we obtained 1,533 functions that meet these criteria.
4.1.2. Benchmark Construction
@@
The manual annotation for each function is resource-intensive and time-consuming. To ensure the efficiency of our experiments, we employed a statistical sampling method [55]. Specifically, we performed sampling on the 1,533 standalone functions with comments and non-empty code. We ensured that the estimated accuracy within the sample remains within a 0.05 error range at a 95% confidence level while ensuring that functions from each code repository are included in the sample. Finally, we acquired 383 functions, which are used for subsequent experiments.
4.2. Metrics
To comprehensively evaluate whether the three awareness are effective and whether the generated code has successfully perceived three awareness, the following metrics are employed to assess the generated code in different aspects:
@@ 4.2. Metrics
  • Precision: This indicates how much of the code generated by the model is reused correctly.
@@
  • Recall: This indicates how much of the code that really needed to be reused was correctly reused by the model.
@@
  • F1 Score:The harmonic mean of Precision and Recall, which measures the overall effectiveness of the model in reusing code.
@@
  • Library Coverage: To evaluate whether the third-party libraries used in the generated code are exist in the code repository’s existing environment, we calculate library coverage. This metric represents the proportion of libraries used in the generated code that are already pre-installed in the code repository’s environment.
@@
  • Lines of Code (LOC): To assess the conciseness of the generated code, we calculate the number of lines of code (LOC). LOC is determined by counting the lines in the code after removing comments. Fewer lines indicate that the generated code is more concise.
  • CodeBLEU [56]: This is an improved version of the BLEU metric, designed for code, and is a weighted average of the lexical, abstract syntax tree, and data flow matches.
@@
Subsequently, we enlist two developers with more than five years of experience in Python programming to conduct independent evaluations. For each generated result, they assess whether the function utilizes local functions, assigning a score of 1 if it does and 0 if it doesn’t. The same process applies to global functions and third-party libraries. Then we calculate the TP, FP, TN, FN, Precision, Recall, F1, and Accuracy for the local functions awareness. After completing their evaluations, they discuss any discrepancies and reach a consensus. After the evaluations, discrepancies are discussed, and a consensus is reached.
In all the upcoming experiments, we use this benchmark for evaluation and utilize the same method to annotate the code generated by the model. The calculated Cohen’s Kappa coefficient [57] for all annotated data across the research questions fall within the range of 86% to 95%, indicating a high level of agreement among the various assessments.
4.3. Foundation Models
As for the LLM selection, we used the gpt-3.5-turbo model, which demonstrates excellent performance in various software engineering tasks. As for the embedding model, we selected the ada-embedding-002 model [58], which has the capability of generating high-quality sentence embeddings, thus enabling us to perform effective similarity computation and retrieval operations. Similar to other LLM-empowered frameworks in SE domain, the more powerful foundation model is selected, the better performance can be achieved, generally.
5. Evaluation
5.1. The Effectiveness of Local-Aware Knowledge (RQ1)

### 03 Mar 2024 v3

@@ III-A2 Third-Party Library Extraction
Third-party libraries often appear in code files, and when specific libraries are provided to a model, it likely has learned how to use them during training [43, 44]. Therefore, supplying just the library name allows the model to automatically understand how to use it for a given task. For example, we can ask the model to use the NumPy library to compute the cosine similarity, given the input vector_a and vector_b. In this case, the model can directly call functions in the NumPy library, such as numpy.dot and numpy.linalg.norm, to perform the required computations. The pre-trained model’s prior knowledge and extensive large-scale training enable them to effectively utilize third-party libraries, improving the efficiency and accuracy of code generation. Therefore, we extract the import statements of third-party libraries directly from all code files via AST and saved their library names to construct the third-party library base.
III-B Three Types of Repo-Aware Knowledge Retrieval
@@
As we discussed in Section II-A, although the maximum input length grows greatly, it is still impossible to feed the entire code repo and third-party library information to LLMs. Furthermore, irrelevant context not only affects model efficiency and increases expenses, but also potentially compromises the model performance [34].
@@
To enable the model to fully utilize the three types of information extracted from the code repository, we conduct comprehensive knowledge mining: extracting four types of local information (Section III-B1), retrieving global functions from the function base using developer requirements to obtain global-aware information (Section III-B2), and structuring third-party library information from the third-party library base (Section III-B3).
III-B1 Local-Module Knowledge Mining
@@
  • Local functions. Local functions encompass all functions that are defined and implemented within the local module. We extract these local functions and represent them in the form of $<$Fully Qualified Name (FQN), Function Summary, Function Signature$>$. The FQN provides details about the function’s location within the module, the Function Summary reflects the natural language description of the function’s purpose, and the Function Signature reveals its input and output specifications.
  • Class Instance Attributes. Class instance attributes (or instance variables) are variables defined within a class’s constructor function (__init__ in Python). These variables are specific to instances of the class and store instance-specific attributes or state information. Some local functions may be defined within a class might involve interactions with class instance attributes, accessing or modifying them to achieve specific functionality.
  • Local Module FQN. Local Module FQN partially reflects the structural information of the working context (local module) within the entire code repository. It helps the model understand the context of how local functions, class instance attributes, and variables within this module are used.
@@
III-B2 Global-Module Knowledge Mining
@@
Figure 3: $A^{3}$ Code Generator Unit
@@
  • What-If Code Generator Unit: To retrieve global functions from the function base, we consider textual and code semantic similarity. For code semantic similarity matching, we utilize the code generation capability of the LLM, inputting user requirements to generate a what-if code, which encapsulates the logic to fulfill those requirements and serves as input for code similarity matching in the Global Function Retrieval Unit. For in-context learning, we provide three representative examples. Notably, the generated code is only used for retrieving the global function, so completeness is not required, as it is not the final code.
@@
  • Global Function Retrieval Unit: This unit uses function descriptions and the what-if function-code as input to retrieve reusable global functions from the function base. Specifically, we first embed function_description into $vector_{des}$ and compute the cosine similarity between this vector and the summary vectors of all functions in the function base. Thus, we can rank and obtain the top k $RetrievedFun_{des}$. Second, we embed the what-if function-code as a $vector_{f}$ and compute the cosine similarity between this vector and the code vectors of all the functions in the function base, so we can rank and obtain the top k $RetrievedFun_{code}$. Finally, we merge the sets of $RetrievedFun_{des}$ and $RetrievedFun_{des}$ to obtain the global function that is most relevant to the provided input, which is denoted as $RetrievedFun$.
@@
  • Information Unification Unit: For each $RetrievedFun$, we organize and display the retrieved information as $<$FQN, function signature, summary$>$ in a structured format.
III-B3 Third-Party Library Knowledge Mining
@@
In the process of code generation, to behavior similar to human developers, it’s crucial for the LLM to be aware of the third-party libraries pre-installed in the current code repository environment. Without this awareness, there’s a risk of generating code that relies on third-party libraries unsupported by the current environment, rendering the code non-functional. To ensure that, we create a comprehensive list of all the third-party libraries found in the third-party library base, as extracted in section III-A2. It’s important to note that we extract only the names of these libraries and do not include any of their code. Third-party libraries are publicly accessible and widely applied, so according to our empirical observation, we found that the model has already encountered and learned most of the libraries during its training process. Thus, we only extract the names of these libraries without any of their code. By constructing this list of third-party libraries, we empower the LLM to generate code that aligns with the available environment and utilizes the appropriate libraries effectively, reducing the risk of non-executable code.
III-C $A^{3}$ode Generation
@@
After section III-B, we obtain structured local-aware knowledge, global-aware knowledge, and third-party-library-aware knowledge. To ensure the model’s full awareness of these three types of knowledge, we design an $\mathbf{A^{3}}$ Code Generation Unit that seamlessly integrates user requirements with the three knowledge sources to generate $A^{3}$ functional code.
The $A^{3}$ Code Generation Unit, as depicted in Fig. 3, consists of four key components: task description, examples, input, and output. In the task description (purple part), we employ a structured prompt to delineate the repository-aware code generation task. Key components of the prompt include: 1) @persona: Guiding the model to assume the role of an expert developer, providing clarity on the task. 2) @terminology: Defining crucial information in a keyword format. 3) @instruction: This section furnishes explicit instructions for executing the task:
  • [leftmargin=*]
  • @command: Offering an overarching strategy for the model to follow during the task.
  • @rule: Outlining the basic limitations and considerations to be observed during code generation.
  • @format: Detailing the expected presentation format for the source code generated for the new function.
@@
In the @command, we guide the model through the code generation process step by step by employing chain-of-thought [45, 46, 47].
@@
  • Additionally, we include an @command that reiterates the importance of strictly following the steps from 1 to 5.
In the @rule, we establish three rules to govern the code generated by the model, including calling import FQN, restricting the use of specified third-party libraries, and ensuring that the generated code is not empty (as the model may occasionally generate abstract functions). We then provide three representative examples. Finally, we input the developer requirement and our 3 knowledge into Unit to get the $A^{3}$ function code (i.e., _to_be_generated_code in Fig. 3 (b)). Notably, developers only needs to input requirements, and all 3 types of repo-aware knowledge are obtained automatically.
Due to the page limitation, we have organized all prompts for our AI units (e.g., Code Summarizer, and $A^{3}$ Code Generator) in the “prompts/” folder of our artifact [19].
IV Experiment Setting
In this section, we outline the default foundation models, dataset, and evaluation metrics for assessing our framework. We introduce four Research Questions (RQs):
@@ IV Experiment Setting
IV-A Foundation Models
We select GPT-3.5-Turbo-16k [48] as the default foundation LLM for this study due to its excellent performance in various SE tasks, along with its popularity and cost-efficiency. As for the embedding model, we use the ada-embedding-002 model [49], which has the capability of generating high-quality sentence embeddings, enabling us to perform effective similarity computation and retrieval operations. Similar to other LLM-empowered frameworks in SE domain, the more powerful model is selected, the better performance can be achieved, generally.
IV-B Dataset Collection
While there exist some benchmark datasets designed for code generation, such as [50, 51, 35, 52, 16], they suffer from the following issues: 1) some of the existing datasets contain code repositories that are already present in the training corpora of popular LLMs, leading to an overestimation of the code repository-aware code generation capabilities. 2) existing benchmarks only evaluate class-level or small unit test cases, which is not suitable for our research goals. 3) Based on our literature review, none of existing code generation benchmark datasets can evaluate whether $a^{3}$ CodGen can effectively obtain, understand, and correctly reuse three distinct types of code repository information. Therefore, to address the potential risk of data leakage and accurately reflect the capability of $A^{3}$- CodGen, we created a new benchmark dataset, dubbed RepoEval. This decision stems from our intention to verify that the model truly grasp $A^{3}$ code generation after receiving information we provide.
IV-B1 Code Repository Collection
According to OpenAI’s official website, the training data for GPT-3.5-Turbo-16k is cut off as of September 2021. To ensure that our benchmark dataset contains code repositories that have not been seen during model training, we select projects with the earliest release date after October 1, 2021 on PyPI. We further filter the GitHub repositories by choosing those with more than 100 stars and 100 branches, as these metrics indicate repository quality. Ultimately, we identify 29 code repositories to construct our benchmark, each averaging 82.517 code files.
Then, we create 29 function bases and 29 third-party library bases for the corresponding code repositories, following the method outlined in Section III-A. The function bases from these 29 code repositories encompass a total of 13,784 functions. Out of these, 10,077 are class member functions, leaving 3,707 as module functions. Since class member functions often involve complex relationships related to inheritance and other classes, it becomes highly labor-intensive to analyze the code invocation relationships to construct the benchmark. Therefore, our benchmark dataset focuses solely on module functions. From these 3,707 module functions, we filter out functions that had comments and were non-empty, and finally, we obtain 1,533 functions.
IV-B2 Benchmark Construction
@@
Manual annotation for each function is resource-intensive and time-consuming. To ensure experiment efficiency, we employ a statistical sampling method [55]. Specifically, we perform sampling on the 1,533 module functions obtained in Section IV-B1. We ensure that the estimated accuracy within the sample remains within a 0.05 error range at a 95% confidence level while ensuring functions from each code repository are included in the sample. Finally, we acquire 383 functions for subsequent experiments.
IV-C Metrics
To comprehensively evaluate whether the three awareness are effective, we employ the following metrics:
@@ IV-C Metrics
  • Reuse Awareness: To assess the model’s ability to correctly reuse local, global, and third-party library information when necessary and avoid reuse when unnecessary, we computed the following awareness metrics:
    • True Positive (TP): The model-generated code correctly reuses local, global, or third-party library information, consistent with the benchmark code. For instance, if the model-generated code reuses local functions and the benchmark code also reuses local functions, it is considered TP in the case of local reuse, as well as in the case of global reuse and third-party library reuse.
    • False Positive (FP): The model-generated code incorrectly reuses local, global, or third-party library information, where no reuse occurs in the benchmark code.
    • False Negative (FN): The model-generated code fails to reuse local, global, or third-party library information, but reuse occurs in the benchmark code.
    • True Negative (TN): The model-generated code correctly does not reuse local, global, or third-party library information, consistent with the benchmark code.
@@
    • Precision: Calculated as $TP/(TP+FP)$, indicating the proportion of the model-generated code that is correctly reused.
@@
    • Recall: Calculated as $TP/(TP+FN)$, indicating the proportion of the code that actually needed to be reused and is correctly reused by the model.
@@
    • F1: The harmonic mean of Precision and Recall, providing an overall measure of the model’s effectiveness in reusing code, calculated as $2*(Precision*Recall)/(Precision+Recall)$.
    • Accuracy: Calculated as $(TP+TN)/(TP+FP+TN+FN)$, representing the percentage of cases where the model is correctly aware or unaware.
  • Reuse Correctness: After identifying local/global/third-party library information, we assess the correctness of reuse. The metric is determined by the number of exact matches between the functions/libraries reused by the model and those in the benchmark, divided by the total number of functions/libraries reused by the model. For instance, in local reuse, if the model reuses functions $local_{a}$, $local_{b}$, and $local_{c}$, and the benchmark reuses functions $local_{b}$, $local_{c}$, and $local_{d}$, there are 2 exact matches, resulting in a correctness score of 2/3 = 0.667.
@@
  • Library Coverage: This metric represents the proportion of libraries used in the generated code that are already pre-installed in the code repository’s environment.
@@
  • Lines of Code (LOC): This metric reflects code conciseness, determined by counting lines after removing comments. Lower LOC indicates more concise code.
@@
Subsequently, we enlist two developers with more than five years of experience in Python programming to conduct independent evaluations. For each generated result, they assess whether the function utilizes local functions, assigning a score of 1 if it does and 0 if it doesn’t. This score contributes to the calculation of reuse awareness. In instances of local function reuse, they identify all the reused functions, forming the basis for calculating reuse correctness. The same process applies to global functions and third-party libraries. After completing their evaluations, they discuss any discrepancies and reach a consensus. In all upcoming experiments, we employ this benchmark for evaluation and annotate the model-generated code using the same method. The Cohen’s Kappa coefficient [56] for annotated data across research questions ranges from 86% to 95%, showing high agreement among assessments.
V Evaluation
V-A The Effectiveness of Local-Aware Knowledge (RQ1)